# Multifidelity Simulation-based Inference for Computationally Expensive Simulators

### v3

@@ 4.4 When does pre-training help?
Providing theoretical guarantees for these questions necessitates a formal characterization of convergence rates in NPE with transfer learning. Although recent works have begun addressing these challenges in NPE (Frazier et al., 2024), current theoretical frameworks of transfer learning (Tahir et al., 2024; Yun et al., 2020; Tripuraneni et al., 2020; Lampinen & Ganguli, 2018), rely on simplifying assumptions (e.g., linear networks) that do not fully capture the complexities of MF-NPE. Given this limitation, we instead empirically explored the conditions in which low-fidelity pre-training facilitates effective transfer learning. To do this, we evaluate MF-NPE where the low- and high-fidelity simulators are related by systematic perturbations (e.g., additive noise or a shift in representation) (Fig. 4). Through these experiments, we identified mutual information and representational similarity as two key dimensions in predicting the effectiveness of pre-training (details in Appendix B).
@@ 5 Discussion
Despite MF-NPE’s advantages, the method comes with some challenges. First, the effectiveness of MF-NPE relies on the similarity between the low-fidelity and high-fidelity models. Fortunately, in many situations, domain experts will know beforehand whether low-fidelity models are poor approximations of high-fidelity models. Second, MF-NPE and MF-TSNPE inherit the limitations of NPE and TSNPE, respectively, in particular regarding the scalability of simulation-based inference to high-dimensional parameter spaces. How to balance exploration of high-dimensional parameter spaces and computational cost in a simulation-based inference setting remains a topic of active research. Third, A-MF-TSNPE requires the training of an ensemble of density estimators, which leads to substantial computational costs in training and hyperparameter tuning. This method should therefore only be preferred in cases where the cost incurred in simulations outweighs the training cost. We estimate this to be the case for the tasks with the multicompartment neuron model and the spiking network model, for which the cost of one simulation and the training of one density estimator are comparable in certain settings (e.g., on the order of minutes, for a network trained on $10^{3}$ samples).
@@ 6 Reproducibility statement
The training and simulation costs for all tasks and SBI methods, as well as a detailed description of the experimental setup, are described in Appendices D.1 and K. The corresponding code and data will be publicly released upon publication.
@@ Acknowledgments
We thank Karthik Sama, Najlaa Mohamed, Guy Moss, Marcel Nonnenmacher and Pierre Vanvolsem for discussions. We also thank the staff of the VIB Data Core for their support. Anastasia N. Krouglova was supported by an FWO grant (G097022N). Hayden R. Johnson was supported by an FWO grant (G053624N). Basile Confavreux was supported by a Schmidt Science Polymath Award to Andrew Saxe, the Sainsbury Wellcome Centre Core Grant from Wellcome (219627/Z/19/Z) and the Gatsby Charitable Foundation (GAT3850). Michael Deistler was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC 2064 – Project number 390727645), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039A) and the European Union (ERC, “DeepCoMechTome”, ref. 101089288). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Michael Deistler is a member of the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
@@ Appendix A Usage of LLMs
Appendix B When does pre-training help?
In this section, we empirically investigate conditions in which low-fidelity pre-training facilitates effective transfer learning in approximating the high-fidelity posterior. Concretely, we study MF-NPE in the special case where the low- and high-fidelity models describe the same stochastic process but differ through a controlled perturbation. This setting lets us vary the amount of shared information between the two simulators in a principled way and quantify the effect on downstream performance.
We hypothesize that the effectiveness of pre-training is associated with two primary factors:
  1. 1.
    Mutual information between the low- and high-fidelity simulators.
  2. 2.
To test this hypothesis, we consider the OU2 process under additive Gaussian noise and/or an information-preserving shift in data representation.
As a shift in data representation, we consider a reversal of the coordinate order of the OU-process output. When $x\in\mathbb{R}^{D}$, we can define this using an anti-diagonal permutation matrix
$T(x)=J\,x,\qquad J_{ij}=\begin{cases}1,&j=D-i+1,\\ 0,&\text{otherwise}.\end{cases}$
which simply reverses the coordinate ordering of $x$. Notably, this transformation is invertible and thus information-preserving.
B.1 Controlled experiment: perturbed OU2 process
In our experiment, we define the following generative process:
  1. 1.
    Sample simulation parameters from the prior: $\mu,\sigma\sim p(\mu,\sigma).$
  2. 2.
    Generate a ”high-fidelity” observation given the sampled parameters: $X_{\text{H}}\sim p\!\left(X\,\middle|\,\mu,\sigma\right)$
  3. 3.
    Draw a noise term: $\varepsilon\sim\mathcal{N}\!\left(0,\,\delta^{2}\right)$
  4. 4.
    Generate a ”low-fidelity” observation: $X_{\text{L}}\sim p\!\left(X\,\middle|\,\mu,\sigma+\varepsilon\right)$
  5. 5.
    Apply the transform: $X_{\text{L}}^{\mathrm{inv}}\;=\;T\!\bigl(X_{\text{L}}\bigr)$, where $T$ is the inversion defined above.
As mutual information is invariant to invertible transformations, we can write
$\mathbb{I}[X_{H};X^{\text{inv}}_{L}]=\mathbb{I}[X_{H};X_{L}]=\mathbb{H}[X_{H}]+\mathbb{H}[X_{L}]-\mathbb{H}[X_{H},X_{L}]\\$ (3)
The quantity $\mathbb{I}[X_{H};X_{L}]$ decreases monotonically with the noise scale $\delta$. At the same time, applying the inversion leaves the information content unchanged while disrupting representational coherence. Fig. 4 illustrates how each manipulation affects the uncertainty coefficient
$\mathbb{UC}[X_{H}|X_{L}]=\frac{\mathbb{I}[X_{H};X_{L}]}{\mathbb{H}[X_{H}]},$
which we estimate empirically using MINE (Belghazi et al., 2018) (Fig. 4B), and MF-NPE performance under a fixed simulation budget of $10^{4}$ low-fidelity simulations and $10^{2}$ high-fidelity simulations (Fig. 4C).
Figure 4: (A) Schematic figure representing lower bound on transfer error ($1/\text{MF-NPE{} performance}$) as a function of mutual information between the low- and high-fidelity models, given a fixed simulation budget. (B) Uncertainty coefficient monotonically decreases with noise parameter $\delta$ and is invariant to data inversion. (C) Empirical results with MF-NPE support the hypothesis that transfer performance is dependent on both mutual information and representational coherence. Note that NPE—with the same high-fidelity simulation budget of $10^{2}$—has similar performance as MF-NPE in the case where the low- and high-fidelity models have low mutual information.
In agreement with our hypothesis, our results suggest the effectiveness of MF-NPE is associated with both the mutual information and representational coherence between low- and high-fidelity simulators (Fig. 4C). Specifically, we find that mutual information is necessary for effective transfer learning, but is not sufficient. Perturbations that preserve information (e.g., invertible transformations) can also critically affect transfer learning performance. Thus, effective pre-training strategies should prioritize high mutual information and coherent representation to maximize downstream performance.
Appendix C Proof of convergence of the NPE log-likelihood loss
@@
Appendix D Further experimental details
D.1 Training procedure
All methods and evaluations were implemented in PyTorch (Paszke et al., 2019). We used the Zuko package (version 1.4.0, MIT License) [^1] https://github.com/probabilists/zuko(Roset, 2024) to implement the normalizing flow, based on the Neural Spline Flows (NSF) architecture (Durkan et al., 2019), and the SBI package (version 0.24.0, Apache 2.0 license) [^2] https://github.com/sbi-dev/sbi (Boelts et al., 2024) for additional functions. The parameters used to generate simulations were logit-transformed for numerical stability, and the summary statistics were z-scored to improve the performance of the normalizing flows. The loss function is the negative-log likelihood, and the optimization function is the Adam optimizer (Kingma & Ba, 2017).
@@ D.1 Training procedure
For the evaluation of A-MF-TSNPE, we used 5 rounds of active sampling, where $80\%$ of the high-fidelity dataset was used for standard MF-NPE training, and $20\%$ was split across the rounds of active sampling. The active samples were selected using the acquisition function over an ensemble of 5 networks.
@@
Appendix E Tasks
E.1 OU process
@@
References (Holý & Tomanová, 2022; Carter & Strey, 2023; Kou et al., 2012)
@@
E.2 Posterior distributions over OU process
@@
E.3 OU process with varying parameter space
@@
Figure 8: MF-NPE benefits from larger low-fidelity datasets. We ran MF-ABC with hyperparameters $\boldsymbol{\epsilon}=(1,1)$ and $\boldsymbol{\eta}=(0.9,0.3)$ (more details in Appendix F.1.1). All variants of our method perform better than MF-ABC and NPE.
E.4 SLCP
Simple Likelihood Complex Posterior (SLCP) is s a benchmark inference task that has been artificially designed to have a simple likelihood, but a very non-trivial 5-dimensional posterior to infer. In this example, we study the impact of multifidelity in cases where the dimensionality of the parameter space differs between the low-fidelity and high-fidelity models.
@@ E.6 Image example
The Gaussian Blob image example contains high-dimensional observations that have been embedded with a CNN embedding from the SBI package (Boelts et al., 2024). The model renders a 2D image, which we modeled as a 256 x 256 pixel image of a Gaussian blob, and aiming to infer three parameters ($\mu_{\text{off}},\sigma_{\text{off}},\gamma$): the horizontal and vertical displacements of the blob, and its contrast (Lueckmann et al., 2019). The image is a grey-scale and is generated through a binomial distribution with a total count of 255 and probability $p_{ij}$, as described in Lueckmann et al. (2019).
@@
Appendix F Gaussian Blob evaluation
@@
F.1 Data generation and transformations for increased network performance
@@
F.1.1 Multifidelity Approximate Bayesian Computation (MF-ABC)
We translated into Python a publicly available Julia implementation of the multifidelity ABC algorithm (Prescott & Baker, 2020). In our setup, the adaptive sampling scheme of MF-ABC selected approximately $30\%$ of the batch size as high-fidelity samples in the OU2 and OU3 tasks, and $50\%$ in the OU4 task. To ensure consistency with our neural network experiments, we z-scored the simulator output before inference. We also explored the effect of varying the acceptance threshold $\epsilon$. We found that the hyperparameters slightly affect the performance of MF-ABC, but that MF-NPE always shows superior performance than MF-ABC (Figure 11). However, MF-ABC has several other hyperparameters to tune. We cannot exclude the hypothesis that larger performance gains could be obtained from such an approach by a more extensive hyperparameter search.
@@ F.1.1 Multifidelity Approximate Bayesian Computation (MF-ABC)
Appendix G Task 2: Multicompartmental single neuron model
The response of a morphologically detailed neuron to an input current is typically modeled with a multicompartmental neuron model wherein the voltage dynamics of each compartment $\mu$ are based on Hodgkin-Huxley equations (Hodgkin & Huxley, 1952):
@@ Appendix G Task 2: Multicompartmental single neuron model
All simulations were performed using Jaxley (V 0.8.2) (Deistler et al., 2024) over 120 ms. The injection current is a step current of 0.55mV over 100 ms, with a delay of 10ms. The step size of the simulator is 0.025.
@@ G.1 NRMSE evaluation
G.2 Simulation-based calibration and Posterior distributions
@@
G.3 Posterior Predictive Checks
@@
G.4 Low and high-fidelity traces
@@
Figure 17: Simulated membrane potential traces of an L5 pyramidal cell (L5PC) model with Jaxley (Deistler et al., 2024). The low- and high-fidelity models are, respectively, a single-compartment model per dendritic branch versus an eight-compartment model per branch.
Appendix H Task 3: Spiking network model
H.1 High-fidelity model
@@
with $w_{ij}(t)$ the connection strength between neurons $i$ and $j$ (unitless), $\delta_{k}(t)=\sum\delta(t-t_{k}^{*})$ the spike train of pre-synaptic neuron $k$, where $t_{k}^{*}$ denotes the spike times of neuron $k$, and $\delta$ the Dirac delta. All neurons received input from 5k Poisson neurons, with 5% random connectivity and constant rate $r_{\text{ext}}=10$Hz in each simulation. The recurrent connectivity was instantiated with random sparse connectivity (10%). All recurrent synapses in the network ($E$-to-$E$ and $E$-to-II, II-to-$E$, II-to-II) underwent variations of spike-timing dependent plasticity (STDP) (Gerstner & Kistler, 2002; Confavreux et al., 2023). Given the learning rate $\eta$, the weights between the neurons $i$ and $j$ of connection type $X$-to-$Y$ evolved over time as:
@@
H.2 Low-fidelity model
Following previous work (Confavreux et al., 2023; Vogels et al., 2011; Dayan & Abbott, 2001), a (partial) mean-field theory applied to the $E$-to-$E$ and $E$-to-II connections in the model described above gave:

### v4

@@ 4.4 When does pre-training help?
Providing theoretical guarantees for these questions necessitates a formal characterization of convergence rates in NPE with transfer learning. Although recent works have begun addressing these challenges in NPE (Frazier et al., 2024), current theoretical frameworks of transfer learning (Tahir et al., 2024; Yun et al., 2020; Tripuraneni et al., 2020; Lampinen and Ganguli, 2018), rely on simplifying assumptions (e.g., linear networks) that do not fully capture the complexities of MF-NPE. Given this limitation, we instead empirically explored the conditions in which low-fidelity pre-training facilitates effective transfer learning. To do this, we evaluated MF-NPE where the low- and high-fidelity simulators were related by systematic perturbations (Fig. 4).
We hypothesized that the effectiveness of pre-training is associated with two primary factors:
  1. 1.
    Mutual information between the low- and high-fidelity simulators.
  2. 2.
To isolate the effects of these factors, we constructed controlled variants of the OU2 process in which the low-fidelity simulator differs from the high-fidelity one through two distinct transformations. In the baseline setup, the simulators generate observations according to
$x\sim p(x\mid\mu,\sigma),\qquad x_{\mathrm{L}}\sim p(x\mid\mu,\sigma+\delta),$
where the perturbation $\delta$ increases the noise of the low-fidelity simulator and therefore reduces $\mathbb{I}[x;x_{\mathrm{L}}]$ monotonically as $\delta$ grows.
Second, to independently manipulate representational coherence, we applied an invertible coordinate-reversal transformation $x_{\mathrm{L}}^{\mathrm{inv}}=T(x_{\mathrm{L}})$, implemented via an anti-diagonal permutation matrix that reverses the ordering of the output dimensions. Because $T$ is invertible, the mutual information between the two simulators is unchanged:
$\mathbb{I}[\boldsymbol{x};\,\boldsymbol{x}_{\mathrm{L}}^{\mathrm{inv}}]\;=\;\mathbb{I}[\boldsymbol{x};\,\boldsymbol{x}_{\mathrm{L}}]\;=\;\mathbb{H}[\boldsymbol{x}]+\mathbb{H}[\boldsymbol{x}_{\mathrm{L}}]-\mathbb{H}[\boldsymbol{x},\boldsymbol{x}_{\mathrm{L}}].$
Thus, while $\mathbb{I}[x;x_{\mathrm{L}}]$ decreases monotonically with the noise scale $\delta$, the inversion leaves the information content unchanged while disrupting representational coherence. Figure 4 illustrates how each manipulation affects the uncertainty coefficient (Figure 4B), which we estimated empirically using MINE (Belghazi et al., 2018), and MF-NPE performance under a fixed simulation budget of $10^{4}$ low-fidelity and $10^{2}$ high-fidelity simulations (Figure 4C).
In agreement with our hypothesis, our results suggest that the effectiveness of MF-NPE depends on both the mutual information and the representational coherence between low- and high-fidelity simulators (Fig. 4C). Specifically, mutual information is necessary for effective transfer learning but not sufficient: perturbations that preserve information (e.g., invertible transformations) can still substantially impair transfer performance. Effective pre-training strategies should therefore prioritize low-fidelity simulators that are both highly informative and representationally aligned with the high-fidelity model.
@@ 5 Discussion
Despite MF-NPE’s advantages, the method comes with some challenges. First, the effectiveness of MF-NPE relies on the similarity between the low-fidelity and high-fidelity models. Fortunately, in many situations, domain experts will know beforehand whether low-fidelity models are poor approximations of high-fidelity models. Second, MF-NPE and MF-TSNPE inherit the limitations of NPE and TSNPE, respectively, in particular regarding the scalability of simulation-based inference to high-dimensional parameter spaces. How to balance exploration of high-dimensional parameter spaces and computational cost in a simulation-based inference setting remains a topic of active research. Third, MF-TSNPE-AF requires the training of an ensemble of density estimators, which leads to substantial computational costs in training and hyperparameter tuning. This method should therefore only be preferred in cases where the cost incurred in simulations outweighs the training cost. We estimate this to be the case for the tasks with the multicompartment neuron model and the spiking network model, for which the cost of one simulation and the training of one density estimator are comparable in certain settings (e.g., on the order of minutes, for a network trained on $10^{3}$ samples).
@@ 6 Reproducibility statement
The training and simulation costs for all tasks and SBI methods, as well as a detailed description of the experimental setup, are described in Appendices C.1 and J. The corresponding code and data are publicly available on Github: github.com/goncalab/multifidelity-NPE.
@@ Acknowledgments
We thank Karthik Sama, Najlaa Mohamed, Guy Moss, Marcel Nonnenmacher and Pierre Vanvolsem for discussions. We also thank the staff of the VIB Data Core for their support. Anastasia N. Krouglova was supported by an FWO grant (G097022N). Hayden R. Johnson was supported by an FWO grant (G053624N). Basile Confavreux was supported by a Schmidt Science Polymath Award to Andrew Saxe, the Sainsbury Wellcome Centre Core Grant from Wellcome (219627/Z/19/Z) and the Gatsby Charitable Foundation (GAT3850). Michael Deistler was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC 2064 – Project number 390727645), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039A) and the European Union (ERC, “DeepCoMechTome”, ref. 101089288; Jakob H. Macke). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Michael Deistler was a member of the International Max Planck Research School for Intelligent Systems (IMPRS-IS). Pedro J. Gonçalves is thankful for the financial support from the Flemish government through long-term structural funding Methusalem (grant METH/26/003).
@@ Appendix A Usage of LLMs
Appendix B Proof of convergence of the NPE log-likelihood loss
@@
Appendix C Further experimental details
C.1 Training procedure
All methods and evaluations were implemented in PyTorch (Paszke et al., 2019). We used the Zuko package (version 1.4.0, MIT License) [^1] https://github.com/probabilists/zuko(Roset, 2024) to implement the normalizing flow, based on the Neural Spline Flows (NSF) architecture (Durkan et al., 2019), and the SBI package (version 0.24.0, Apache 2.0 license) [^2] https://github.com/sbi-dev/sbi (Boelts et al., 2024) for additional functions. The parameters used to generate simulations were logit-transformed for numerical stability, and the summary statistics were z-scored to improve the performance of the normalizing flows. The loss function is the negative-log likelihood, and the optimization function is the Adam optimizer (Kingma and Ba, 2017).
@@ C.1 Training procedure
For the evaluation of MF-TSNPE-AF, we used 5 rounds of active sampling, where $80\%$ of the high-fidelity dataset was used for standard MF-NPE training, and $20\%$ was split across the rounds of active sampling. The active samples were selected using the acquisition function over an ensemble of 5 networks.
@@
Appendix D Tasks
D.1 OU process
@@
References (Holý and Tomanová, 2022; Carter and Strey, 2023; Kou et al., 2012)
@@
D.2 Posterior distributions over OU process
@@
D.3 OU process with varying parameter space
@@
Figure 8: MF-NPE benefits from larger low-fidelity datasets. We ran MF-ABC with hyperparameters $\boldsymbol{\epsilon}=(1,1)$ and $\boldsymbol{\eta}=(0.9,0.3)$ (more details in Appendix E.1.1). All variants of our method perform better than MF-ABC and NPE.
D.4 Inferring the parameters of a Gaussian model pretrained on the OU3 model
In this example, we examine how the performance changes when the low-fidelity model has a larger number of parameters than the high-fidelity model: the low-fidelity model is the Ornstein-Uhlenbeck process with three parameters, and the high-fidelity model corresponds to i.i.d. Gaussian samples parameterised by a mean and variance (so, only two parameters). To accomplish that, the density estimator pre-trained on the low-fidelity model was fine-tuned only on the dimensions of the high-fidelity and the extra dimension was kept as a dummy dimension. NPE was directly trained on the 2-dimensional parameter space of the high-fidelity model. At inference time, the posterior evaluation was performed only on the high-fidelity parameter dimensions. We observe that when the dimension of $\theta$ is smaller than the dimension of $\theta_{L}$, transfer learning provides a significant improvement in performance.
Figure 9: Evaluation with C2ST and MMD over a two-dimensional Gaussian Samples model, pretrained on the three-dimensional OU process model.
D.5 SLCP
Simple Likelihood Complex Posterior (SLCP) is a benchmark inference task that has been artificially designed to have a simple likelihood, but a very non-trivial 5-dimensional posterior to infer. In this example, we study the impact of multifidelity in cases where the dimensionality of the parameter space differs between the low-fidelity and high-fidelity models.
@@ D.7 Image example
The Gaussian Blob image example contains high-dimensional observations that have been embedded with a CNN embedding from the SBI package (Boelts et al., 2024). The model renders a 2D image, which we modeled as a 256 x 256 pixel image of a Gaussian blob, and aiming to infer three parameters ($\mu_{\text{off}},\sigma_{\text{off}},\gamma$): the horizontal and vertical displacements of the blob, and its contrast (Lueckmann et al., 2019). The image is in grey-scale and is generated through a binomial distribution with a total count of 255 and probability $p_{ij}$, as described in Lueckmann et al. (2019).
@@
Appendix E Gaussian Blob evaluation
@@
E.1 Data generation and transformations for increased network performance
@@
E.1.1 Multifidelity Approximate Bayesian Computation (MF-ABC)
We translated into Python a publicly available Julia implementation of the multifidelity ABC algorithm (Prescott and Baker, 2020). In our setup, the adaptive sampling scheme of MF-ABC selected approximately $30\%$ of the batch size as high-fidelity samples in the OU2 and OU3 tasks, and $50\%$ in the OU4 task. To ensure consistency with our neural network experiments, we z-scored the simulator output before inference. We also explored the effect of varying the acceptance threshold $\epsilon$. We found that the hyperparameters slightly affect the performance of MF-ABC, but that MF-NPE always shows superior performance than MF-ABC (Figure 12). However, MF-ABC has several other hyperparameters to tune. We cannot exclude the hypothesis that larger performance gains could be obtained from such an approach by a more extensive hyperparameter search.
@@ E.1.1 Multifidelity Approximate Bayesian Computation (MF-ABC)
Appendix F Task 2: Multicompartmental single neuron model
The response of a morphologically detailed neuron to an input current is typically modeled with a multicompartmental neuron model wherein the voltage dynamics of each compartment $\mu$ are based on Hodgkin-Huxley equations (Hodgkin and Huxley, 1952):
@@ Appendix F Task 2: Multicompartmental single neuron model
All simulations were performed using Jaxley (V 0.8.2) (Deistler et al., 2025) over 120 ms. The injection current is a step current of 0.55mV over 100 ms, with a delay of 10ms. The step size of the simulator is 0.025.
@@ F.1 NRMSE evaluation
F.2 Simulation-based calibration and Posterior distributions
@@
F.3 Posterior Predictive Checks
@@
F.4 Low and high-fidelity traces
@@
Figure 18: Simulated membrane potential traces of an L5 pyramidal cell (L5PC) model with Jaxley (Deistler et al., 2025). The low- and high-fidelity models are, respectively, a single-compartment model per dendritic branch versus an eight-compartment model per branch.
Appendix G Task 3: Spiking network model
G.1 High-fidelity model
@@
with $w_{ij}(t)$ the connection strength between neurons $i$ and $j$ (unitless), $\delta_{k}(t)=\sum\delta(t-t_{k}^{*})$ the spike train of pre-synaptic neuron $k$, where $t_{k}^{*}$ denotes the spike times of neuron $k$, and $\delta$ the Dirac delta. All neurons received input from 5k Poisson neurons, with 5% random connectivity and constant rate $r_{\text{ext}}=10$Hz in each simulation. The recurrent connectivity was instantiated with random sparse connectivity (10%). All recurrent synapses in the network ($E$-to-$E$ and $E$-to-II, II-to-$E$, II-to-II) underwent variations of spike-timing dependent plasticity (STDP) (Gerstner and Kistler, 2002; Confavreux et al., 2023). Given the learning rate $\eta$, the weights between the neurons $i$ and $j$ of connection type $X$-to-$Y$ evolved over time as:
@@
G.2 Low-fidelity model
Following previous work (Confavreux et al., 2023; Vogels et al., 2011; Dayan and Abbott, 2001), a (partial) mean-field theory applied to the $E$-to-$E$ and $E$-to-II connections in the model described above gave: